In the aggregate expenditure model for a closed economy, assuming investment, government spending and taxes are exogenous,if the marginal propensity to consume is 0.8 , simultaneou 50 unit increase in government spending and a 20 unit decrease in investment will change equilibrium income by



i.87.5 units
ii.350 units
iii.500 units
iv.150 units
1
Expert's answer
2021-09-25T12:32:04-0400

Let the equilibrium income is Y.

Equilibrium income must equal to aggregate expenditure (AE), which is the aggregate of consumption spending (C), Investment spending (I), and government spending (G).

Therefore, AE = C + I + G

Now, by the condition,

Y = AE

Y = C + I + G

Y = 0.8Y + (- 20) + 50

Y – 0.8Y = - 20 + 50

0.2Y = 30

"Y = \frac{30 }{ 0.2} = 150"

Answer: iv.150 units